Grade 9Mechanics

a body of mass 2kg makes an elastic collision with another body at rest and continues to move in the original direction with one fourth of its original speed. the mass of the second body which collides with the first body is
​1) 2kg
2) 1.2kg
3) 3kg
4) 1.5kg

gurnoor kaur
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
Using formula for elastic collision when the target (i.e. second body here) is at rest before collision. 

v1 = (m1 - m2)/(m1 + m2) * u1 
It is given v1 = u1/4 and substituting the value m1 = 2.0 kg, 
u1/4 = (2 - m2)/(2 + m2) * u1 
Divide by u1 
1/4 = (2 - m2)/(2 + m2) 
Or 2 + m2 = 4 * (2 - m2) 
Or 2 + m2 = 8 - 4m2 
Or m2 + 4m2 = 8 - 2 
Or 5m2 = 6 
Or m2 = 6/5 = 1.2 kg 
Ans: 1.2 kg